a) Solve the differential equation
dy/dx + xy= y^2 e^(x^2/2) sinx

b) Given that
y1(x) = x^(-1) is one solution of the differential equation

2x^2 y ′′ + 3xy′ −y =0 ,x >0,
find a second linearly independent solution of the equation
1
Expert's answer
2020-04-03T14:24:59-0400

a)

"\frac{dy}{dx}+xy=y^2e^{\frac{x^2}{2}}\sin x\\
y'+xy=y^2e^{\frac{x^2}{2}}\sin x"

We divide by "y^2"

"\frac{y'}{y^2}+\frac{x}{y}=e^{\frac{x^2}{2}}\sin x\\"

Replacement

"y^{-1}=z, y=z^{-1}, y'=-z^2\cdot z'\\
-\frac{z^{-2}z'}{z^{-2}}+xz=e^{\frac{x^2}{2}}\sin x\\
-z'+xz=e^{\frac{x^2}{2}}\sin x\\
i) -z'+xz=0\\
\frac{dz}{dx}=xz\\
\frac{dz}{z}=xdx\\
ln|z|=\frac{x^2}{2}+ln|c|\\
z_0=c\cdot e^{\frac{x^2}{2}}\\
ii) z=c(x)\cdot e^{\frac{x^2}{2}}\\"

Input "z" in the equation "-z'+xz=e^{\frac{x^2}{2}}\sin x"

"-c'e^{\frac{x^2}{2}}-xe^{\frac{x^2}{2}}c+xe^{\frac{x^2}{2}}c=e^{\frac{x^2}{2}}\sin x\\
c'=-\sin x\\
c=\cos x+c_1"

then "z=(\cos x+c_1)e^{\frac{x^2}{2}}"

or "y^{-1}=(\cos x+c_1)e^{\frac{x^2}{2}}"

and solutoin of equation is function

"y=0"



b)

"y_1(x)=\frac{1}{x}\\
2x^2y''+3xy'-y=0"

for equation

"a_0(x)y''+a_1(x)y'+a_2(x)y=0"

used formula Ostrohrad-Liouville

"\begin{vmatrix}
 y_1 & y_2 \\
 y'_1 & y'_2
\end{vmatrix}=ce^{-\int \frac{a_1(x)}{a_0(x)}dx}"

in our case

"a_0(x)=2x^2, a_1(x)=3x, a_2(x)=-1"

then

"\begin{vmatrix}
 y_1 & y_2 \\
 y'_1 & y'_2
\end{vmatrix}=ce^{-\int \frac{3x}{2x^2}dx}\\
y_1y'_2-y'_1y_2=ce^{- \frac{3}{2}\int\frac{1}{x}dx}\\
y_1y'_2-y'_1y_2=ce^{- \frac{3}{2}ln|x|}\\
y_1y'_2-y'_1y_2=cx^{- \frac{3}{2}}"

Divide by "y_1^2"

"\frac{y_1y'_2-y'_1y_2}{y_1^2}=cx^{- \frac{3}{2}}x^2\\
(\frac{y_2}{y_1})'=cx^{\frac{1}{2}}\\
\frac{y_2}{y_1}=\int cx^{\frac{1}{2}}dx\\
\frac{y_2}{y_1}=\frac{2}{3}cx^{\frac{3}{2}}+c_1"

Then the solution of equation is

"y=(\frac{2}{3}cx^{\frac{3}{2}}+c_1)\frac{1}{x}"

or

"y=\frac{2}{3}cx^{\frac{1}{2}}+c_1\frac{1}{x}"

So second linearly independent solution of the equation is

"y_2=\frac{2}{3}x^{\frac{1}{2}}"